### v1

# Mean field model of contagion processes in urban traffic networks

Abstract
Theoretical arguments and empirical evidence for the emergence of macroscopic epidemic type behavior, in the form of Susceptible-Infected-Susceptible (SIS) or Susceptible-Infected-Recovered (SIR) processes in urban traffic congestion from microscopic network flows is given. Moreover, it’s shown that the emergence of SIS/SIR implies a relationship between traffic flow and density, which is consistent with observations of the so called Fundamental Diagram of Traffic, which is a characteristic signature of vehicle movement phenomena that spans multiple scales. Our results provide a plausible explanation for this scale-spanning signature and put in more firm grounds recent findings that indicate that traffic congestion at the aggregate level can be modeled by simple contagion dynamics.
1 Introduction
Vehicular movement in urban networks can be understood like an out of equilibrium many-body system with a hierarchy of description levels and relevant scales. From a microscopic standpoint, traffic is a granular flow with strong interactions among particles [1]. These interactions are in part a consequence of decision making of the individual particles themselves, which in this sense can be regarded as agents whom can display non-linear and stochastic behavior. At a macroscopic level on the other hand, traffic can be regarded as a continuous flow in terms of coarse grained observables like densities and average velocities, being perhaps manageable by means of interventions on road geometry and network topology intended to reduce traffic jams [2, 3, 4]. Urban traffic capacity is characterized by a Fundamental Diagram of Traffic (FDT), which indicates the existence of a free-flow phase and a congested phase with a transition between them in a critical vehicle density point. This characteristic FDT has been observed to span several scales ranging from the individual street level up to aggregate dynamics over the urban network. At the largest urban scales, traffic congestion has been shown to display a remarkable empirical correspondence with epidemics and propagation by contagion processes [5, 6, 7]. In the present work we put forward a possible unifying framework that explains the scale spanning FDT and the epidemic large scale dynamics by identifying them as emergent properties of an out of equilibrium system that shares commonalities with both granular and continuous flows. The vicinity of road intersections are unavoidable sources of discontinuity and therefore, the places at which the granular nature of the flow becomes more apparent. In other regions far away from the intersections, the continuous density description dominates if additional sources of instability, like bottlenecks, are absent. Therefore we propose a microscopic contact at the crossroads mechanism for the transmission of flow in a transport network. Our resulting model, detailed in the next section, is conceptually similar to the point queue and traffic flow models (see for instance [8] for a discussion on these approaches) with the crucial difference that the transmission rate parameters are not defined by local characteristics of the intersection. In our contagion at junctions model, the rate parameters are instead determined by the total transport in a given time period and the average speed in the road that receives the vehicles. The rate parameters are therefore a consequence of the local discontinuity produced at junctions where incoming vehicles must adapt to the average flow of the receiving road, being this average flow determined by non local spatial and temporal constraints.
2 Crossroad contagion network traffic flow model
Urban traffic has been shown to be consistent with an out of equilibrium granular flow process [1]. Consider therefore a network of links (edges), each representing idealized roads through which such a flow occurs. It’s additionally assumed that aggregates of cars traversing a link $n$ can be described in terms of a continuous car density flow at any point of the link except at crossroads, which can be regarded as contact points in which vehicles enter or leave the $n$-th link. The master equation, or instantaneous transition time limit of the Chapman-Kolmogorov equation [9], for the car density at an edge $n$ is given by,
$\frac{d\rho_{n}(t)}{dt}=\sum_{m\neq n}q_{t}(n|m)\rho_{m}(t)-\sum_{r\neq n}q_{t}(r|n)\rho_{n}(t)$ (1)
The $q_{t}(r|l)$ terms in Eq.(1) represent the instantaneous flow rate of cars from an edge $l$ to another directly connected edge $r$. We now show how the general flow traffic formulation given by Eq.(1), leads naturally to a contagion type model. Consider an intersection in which a link $l$ sends its vehicles to a directly connected link $r$. Assume that there is no other link connected to $r$ besides $l$ in a neighborhood at a distance $\delta x_{r}$ from the intersection $(r,l)$ (see the Figure 1). Denote by $\beta_{r}\equiv\frac{\delta x_{r}}{\delta t}$, the instantaneous average car speed at the interval $\delta x_{r}$ of the $r$-th edge. The number of vehicles $\delta c_{r,l}(t)$ that enter $r$ from $l$ during a time interval $(t,t+\delta t)$ can be expressed equivalently in terms of the flux from $l$ to $r$, denoted by $q(r|l)$, or in terms of the average transiting velocity in the intersection neighborhood $\delta x_{r}$,
$\delta c_{r,l}(t)=q(r|l)\delta t=\frac{q(r|l)\delta x_{r}}{\beta_{r}}$ (2)
Therefore,
$q_{t}(r|l)=\beta_{r}\frac{\delta c_{r,l}}{\delta x_{r}}$ (3)
Now let’s introduce the additional assumption that at the vicinity of the intersection $(r,l)$, the vehicle density of the $r$-th link, $\rho_{r}$ is given by $\rho_{r}=\frac{\delta c_{r,l}}{\delta x_{r}}$. This assumption simply reflects the model setup in which there is no vehicle source for the road $r$ other than road $l$ in the neighborhood $\delta x_{r}$. Equation (3) then leads to the relation,
$q_{t}(r|l)=\beta_{r}\rho_{r}$ (4)
From equation (4), the vehicle density increment $\delta\rho_{r}$ in the time interval $\delta t$, in the road $r$ due to the incoming flow from the intersecting road $l$ is given by
$\delta\rho_{r}=\beta_{r}\rho_{r}\rho_{l}\delta t$ (5)
Now suppose that the $r$-th edge has another crossroad by which it’s directly connected to a link $k$ that acts like the venting point for its flow. By the same arguments that lead to the equation (5), the vehicle density change in the road $r$ due to its intersections with $l$ and $k$ is now given by,
$\frac{\delta\rho_{r}}{\delta t}=\beta_{r}\rho_{r}\rho_{l}-\beta_{k}\rho_{k}\rho_{r}$ (6)
By the equation (6), the master equation (1) can be written in terms of the adjacency matrix of directed links that send flow towards a given link $i$, $\mathcal{A}(i|j)$ as,
$\frac{d\rho_{n}(t)}{dt}=\sum_{m\neq n}\mathcal{A}(n|m)\beta_{n}\rho_{n}\rho_{m}-\sum_{k\neq n}\mathcal{A}(k|n)\beta_{k}\rho_{k}\rho_{n}$ (7)
$=\rho_{n}\left(\beta_{n}\sum_{m\neq n}\mathcal{A}(n|m)\rho_{m}-\sum_{k\neq n}\mathcal{A}(k|n)\beta_{k}\rho_{k}\right)$
Note that the second term inside the parenthesis of Eq.(7) acts like a venting rate for the edge $n$, call it $\gamma_{n}$. Regarding the first term inside of the parenthesis, it can be rewritten by the use of the identity $\sum_{m}\mathcal{A}(n|m)\rho_{m}=\rho_{n}+\sum_{m\neq n}\mathcal{A}(n|m)\rho_{m}$. By the given considerations, the equation (7) therefore reads,
$\frac{d\rho_{n}(t)}{dt}=\rho_{n}\left[\beta_{n}\left(\langle\rho(t)\rangle-\rho_{n}\right)-\gamma_{n}\right]$ (8)
In the previous equation, the quantity $\langle\rho(t)\rangle$ represents the average density of vehicles inside the sub graph formed by $M$ directed links that send vehicles to $n$ together with the $n$-th link itself,
$\langle\rho(t)\rangle\equiv\sum_{m=1}^{M}\mathcal{A}(n|m)\rho_{m}(t)\equiv\bar{\rho}$ (9)
The equation 9 should be interpreted as the average density traversing the link $n$ arrived from its directly connected roads in a time interval $\delta t$. Under an homogeneous mean field description,
$\lim_{M\to\infty}\sum_{m=1}^{M}\mathcal{A}(n|m)\rho_{m}(t)=\bar{\rho}$ (10)
The homogeneous mean field vehicle density dynamics on the $n$-th road is therefore governed by,
$\frac{d\rho_{n}}{dt}=\rho_{n}\left[\beta_{n}(\bar{\rho}-\rho_{n})-\gamma_{n}\right]$ (11)
[…]
$\frac{d\rho_{n}}{dt}=\rho_{n}\left[(1-\rho_{n})\beta_{n}-\gamma_{n}\right]\quad\text{subject to}$ (12)
[…]
$\rho_{n}(\infty)=0,\quad 0\leq\rho_{n}(t)\leq 1\quad\forall\quad n,t.$
The constraints in Eq. (12) model the distinctiveness of each $24$ hour time span due to the weekly seasonality. These constraints can be satisfied by introducing additional variables defined through,
$I\equiv N_{n}-S-R,$ (13)
[…]
In the above equations, $N_{n}$ is the total transport on link $n$ during a given $24$ hour time span. The definitions given by Eq.(13) imply that,
$\frac{dS}{dt}+\frac{dI}{dt}+\frac{dR}{dt}=0,$ (14)
$s_{n}(t)+\rho_{n}(t)+r_{n}(t)=C.$
By choosing $C=1$ without loss of generality, the constrained Eq. (12) directly leads to a Susceptible-Infected-Recovered (SIR) type model for the traffic flow over the $n$-th node, in which each daily shift can be interpreted as an individual contagion process,
$\dot{s}_{n}(t)$ $=$ $-\beta_{n}\rho_{n}(t)s_{n}(t)$ (15)
$\noindent\dot{\rho}_{n}(t)$ $=$ $\beta_{n}\rho_{n}(t)s_{n}(t)-\gamma_{n}\rho_{n}(t)$
$\noindent\dot{r}_{n}(t)$ $=$ $\gamma_{n}\rho_{n}(t).$
In the above equations, $s_{n}$ can be interpreted as the instantaneous density of vehicles in the sub graph conformed by the incoming directly connected links to $n$ and $r_{n}$ the instantaneous density of vehicles in the sub graph conformed by the directly connected links outwards from $n$. The constant flow setup equation (11), is on the other hand equivalent to a SIS contact process in which $\rho_{n}$ is interpreted as the probability that the edge $n$ becomes infected and $1-\rho_{n}$ is the probability of the edge $n$ being susceptible. The SIS/SIR descriptions can be expected to hold in a macroscopic limit under vanishing fluctuations. Define $\beta_{M}=\frac{1}{M}\sum_{m=1}^{M}\beta_{m}$, the average transiting speed over an entire network with $M$ roads. Define also, $\gamma_{M}=\frac{1}{M}\sum_{m=1}^{M}\gamma_{m}$. The average of the dynamics over the graph $\frac{d\rho_{M}}{dt}=\frac{1}{M}\sum_{m}\frac{d\rho_{m}}{dt}$, then reads,
$s_{M}(t)$ $=$ $s_{M}(0)e^{-\beta_{M}\rho_{M}(t)}$ (16)
$\frac{d\rho_{M}(t)}{dt}$ $=$ $\beta_{M}\rho_{M}(t)s_{M}(t)-\gamma_{M}\rho_{M}(t)+\left[\left\langle\left(\frac{d\rho_{M}}{dt}\right)^{2}\right\rangle-\left\langle\left(\frac{d\rho_{M}}{dt}\right)\right\rangle^{2}\right]$
Under vanishing fluctuations at large $M$, the average dynamics over the network converges to the classical compartmental SIR process.
Figure 1: Relation for the the total number of vehicles traversing the $n$-th link with respect to the instantaneous share of vehicles given by its incoming connected roads and expelled to its outwardly directed roads. The contagion at the crossroads model for traffic proposes that the incoming vehicles should adapt to the average speed of the receiving link, which make the crossroads contact points lasting sources of instability.
2.1 Comparison with related approaches and the Fundamental Diagram of Traffic
Descriptions of traffic based on contagion dynamics models have been previously proposed, but these depend on ad-hoc definitions of congested states. For instance, in the seminal work [5], congestion is defined by mapping to bond percolation on a small world network, where the congested states can be transmitted according to a contagion rate parameter and a percolation probability parameter. More recently in [6], congested states are formalized in terms of a threshold for the instantaneous average speed on a link relative to its maximum speed. Given this definition for a congested state road, is then proposed in [6] that congestion spreads through the network by a mechanism that at the aggregated level is described by a simple compartmental SIR dynamics. Ample empirical evidence is given in [6], that indicates that the compartmental SIR behavior closely match the temporal evolution of congestion in a given city.
In our model, congestion is not treated like a state of a link or group of links, but rather emerges in the form of a shape for flow vs density given a total transport in a given time. From Eq. (16) it follows that (with $s_{0}\equiv s(0)$), $\dot{\rho}(\rho)$ is expressed by,
$\dot{\rho}(\rho)\approx\beta\rho s_{0}e^{-\beta\rho}-\gamma\rho$ (17)
The resulting shape is consistent with the Fundamental Diagram of Traffic (FDT) [8], in the sense that it permits to distinguish between free flow and congested phases by the sign of the derivative of the flow with respect to the density. See the upper panel of the Figure 2 for an empirical estimation of an FDT at a local single detector mean field level.
From Eq. (17), it follows that the macroscopic critical density point is approximately given by the solution of the transcendental equation,
$\rho_{c}=1-\frac{\gamma}{\beta s_{0}}e^{\beta\rho_{c}}$ (18)
3 Comparisons with empirical data
Our contagion network traffic flow model has been tested on representative instances of the publicly available UTD19 database [4]. This database provides fundamental traffic variables including flow and density, from 41 cities worldwide. The data is collected from stationary traffic sensors, most of them inductive loop detectors, which report traffic flow and the fraction of time during an observation period in which a detector is occupied as a proxy for the vehicle density.
The network contagion model predicts a SIR dynamics governing the vehicle flow over $24$ hour periods. The upper and lower panels of the Figure 2, show the fitting of the SIR density dynamics from a single detector over a particular street, in the Aussersihl district of Zurich, Switzerland. This instance of the UTD19 database has been showcased in a number of works and has good quality sensor data [4]. Fitting of the SIR rate parameters is done by minimizing the total squared error between the observed densities and those predicted by the SIR dynamics integrated by first order discretization through the Euler method. Because the street density data is a proxy, additional parameters for a linear transformation that resolves the particular sensor calibration are introduced. The resulting optimization problem has two levels. Firstly the rate parameters are estimated by solving
$\textit{min}_{\beta_{n},\gamma_{n}}\quad\sum_{t=1}^{T}[\hat{\rho}_{n}(\beta_{n},\gamma_{n},t)-\rho_{n}(t)]^{2}$ (19)
In Eq.(19), $\hat{\rho}_{n}$ is the vehicle density in the $n$-th link given by the one step ahead prediction of the SIR model at the observed time $t$. The sum of squared errors is calculated from the total number of observations $T$, over the $24$ hour period. The local fitting of the SIR dynamics has been done using the Broyden–Fletcher–Goldfarb–Shanno algorithm, assuming time independent rate parameters, $\beta_{n}(t)=\beta_{n}$, $\gamma_{n}(t)=\gamma_{n}$ and also assuming that $s(0)=1$.
In a second stage, optimal calibration parameters $s_{0},c_{1},c_{2}$ are calculated by,
$\textit{min}_{s_{0},c_{1},c_{2}}\quad\sum_{t=1}^{T}\left[\hat{y}_{n}(s_{0},c_{1},c_{2},t)-\frac{d\rho_{n}(t)}{dt}\right]^{2}$ (20)
$\hat{y}_{n}(s_{0},c_{1},c_{2},t)\equiv s_{0}\beta_{n}\rho_{n}(t)e^{-\beta_{n}\rho_{n}(t)}-\gamma_{n}\rho_{n}(t)$
The calibration parameter $10<s_{0}<$ 1 estimates the fraction of the overall transport covered by the $n$-th link in the mean field limit. The parameters $c_{1}$ and $c_{2}$ are mere calibration constants for the rates, $\beta_{n}\to c_{1}\beta_{n}$, $\gamma_{n}\to c_{2}\gamma_{n}$.
The upper panel of the Figure 2 displays the estimated and observed FDT, $\hat{y}_{n}(\rho_{n})$ and $\dot{\rho}_{n}(\rho_{n})$ respectively. The lower panel on the other hand, shows the estimated and observed SIR temporal dynamics of the vehicle density. Figure 3 reports the comparison between the predicted and observed total transport in the $24$ hour period for an arbitrary set of detectors from links in the vicinity and same district of the detector reported in the Figure 3. The links rate parameters are estimated by the previously discussed procedure individually. The estimated rate parameters averaged over the group of links are consistent with what expected from the equation (16).
Figure 2: Traffic dynamics from a single detector of the Schimmelstrasse street in the Aussersihl district of Zurich, Switzerland, at Wednesday 28th October, 2015. Empirical data is shown in black while the fit of the local mean field contagion model to the data is shown in red.
Figure 3: Predicted (red triangles) and observed (black circles) transport in 22 detectors from the Aussersihl district in Zurich, Switzerland, at Wednesday 28th October of 2015. From the estimated transmission rate parameters, the mean field crossroad contagion network traffic flow model predicts an average road velocity over the covered region of $30.92$ km/hr and an average venting rate of $6.54$ km/hr. The arbitraily chosen detectors from the UTD19 database are, K14D15, K14D13, K14D14, K17D13, K17D14, K17D15, K17D12, K17D11, K13D14, K13D13, K13D12, K14D12, K11D19, K11D11, K12D16, K133D11, K133D12, K133D13, K11D12, K11D14, K12D14, K12D17.
4 Discussion
Despite being an idealization, the contagion framework to traffic appears to capture features that are essential to characterize urban transport at different spatial and temporal scales. The formalism integrates granular with continuous aspects of vehicle flow in conjunction with basic temporal constraints, leading to a local mean field contagion type model. Our empirical comparisons indicate that single detector data aggregates information from all of the links directly connected to the detector’s road. The macroscopic mean field equation (16), predicts that aggregating links in a given macroscopic area of the roads network should converge to macroscopic rate parameters that reflect the vehicle flow in the corresponding network’s section. This appears to be consistent with the limited empirical study reported in the Figure 3. A more thorough experimental study is currently under development by the authors of the present article.
5 Conclusions and future work
In this work it has been introduced a mean field model for vehicle flow in traffic networks that takes into account granular features of the flow at the link’s junctions. This is accomplished by a SIR-type contagion dynamics that aggregates the vehicles sent to a given link. The link level rate parameters are determined by the total transport that traverse the link in a given time interval. We have at this point provided comparisons with empirical data that support our model at individual links and in groups of not necessarily directly connected links inside a macroscopic zone of the network. Clearly a next important step is to apply our approach to large macroscopic traffic networks datasets, by carefully taking into account the network topology. Our equation (18) indicates that the free flow optimal state of a macroscopic traffic network occurs in a critical density given in terms of aggregated rate parameters $\beta$, $\gamma$ and a normalization parameter associated with the total transport in the given time period, $s_{0}$.
Generalizations to the mean field setup that consider network’s spatiotemporal structures in greater detail, can in principle be explored by taking advantage of recent developments in the study of contagion models on networks with complex topologies, like for instance those presented in [10].
6 Acknowledgments
The authors acknowledge to the creators and maintainers of the UTD19 database. The UTD19 database is hosted in the site utd19.ethz.ch