### v3

# Contagion mean field model for transport in urban traffic networks

Abstract
Theoretical arguments and empirical evidence for the emergence of macroscopic epidemic type behavior, in the form of Susceptible-Infected-Susceptible (SIS) or Susceptible-Infected-Recovered (SIR) processes in urban traffic congestion from microscopic network flows is given. Moreover, it’s shown that the emergence of SIS/SIR implies a relationship between traffic flow and density, which is consistent with observations of the so called Fundamental Diagram of Traffic (FDT), which is a characteristic signature of vehicle movement phenomena that spans multiple scales. Our results put in more firm grounds recent findings that indicate that traffic congestion at the aggregate level can be modeled by simple contagion dynamics.
1 Introduction
Vehicular movement in urban networks can be understood like an out of equilibrium many-body system with a hierarchy of description levels and relevant scales. From a microscopic standpoint, traffic is a granular flow with strong interactions among particles [nobottle]. These interactions are in part a consequence of decision making of the individual particles themselves, which in this sense can be regarded as agents which can display non-linear and stochastic behavior. At a macroscopic level on the other hand, traffic can be regarded as a continuous flow in terms of coarse grained observables like densities and average velocities, being perhaps manageable by means of interventions on road geometry and network topology intended to reduce traffic jams [cassidy, dganzo, loder]. Urban traffic capacity is characterized by a Fundamental Diagram of Traffic (FDT), which indicates the existence of a free-flow phase and a congested phase with a transition between them in a critical vehicle density point, as further illustrated in the simple triangular version of the flow vs density FDT schematics provided in the Figure 1 This characteristic FDT has been observed to span several scales ranging from the individual street level up to aggregate dynamics over the urban network. At the largest urban scales, traffic congestion has been shown to display a remarkable empirical correspondence with epidemics and contagion propagation processes, by fitting observed congestion to compartmental epidemic models [saberi, duan]. In the present work we put forward a possible unifying framework that explains the scale spanning FDT and the epidemic large scale dynamics by identifying them as emergent properties of an out of equilibrium system that shares commonalities with both granular and continuous flows. The vicinity of road intersections are unavoidable sources of discontinuity and therefore, the places at which the granular nature of the flow becomes more apparent. In other regions far away from the intersections, the continuous density description dominates if additional sources of instability, like bottlenecks, are absent. Therefore we propose a microscopic contact at the crossroads mechanism for the transmission of flow in a transport network. Our resulting model, detailed in the next section, is conceptually similar to the point queue and traffic flow models (see for instance [blubook-vol1-v10] for a discussion on these approaches) with the crucial difference that the transmission rate parameters are not defined by local characteristics of roads and junctions. Moreover, in contrast to previous approaches that use epidemic models as a metaphor for traffic modeling [sun, saberi, duan], our proposed formalism introduced in Section 2, shows that the mathematical conservation of flux at intersections naturally yields the FDT shape. Empirical evidence for this result is provided in Section 3. Our first principles derivation of the Fundamental Diagram of Traffic has theoretical interest on its own and also potential applications for traffic control, as discussed in Sections 2 and 4. By identifying traffic congestion as mathematically equivalent to a transmitted disease, our model gives possible immunization strategies.
Figure 1: Schematic representation of the Fundamental Diagram of Traffic (FDT). The triangular shape illustrates the two distinct regimes: the free-flow phase (blue solid line) where flow increases linearly with density, and the congested phase (red dashed line) where interactions reduce flow as density approaches the jam density ($\rho_{jam}$). The peak represents the road capacity ($q_{max}$) at the critical density ($\rho_{c}$).
2 Mean field network traffic flow model of contagion at the crossroads
Figure 2: Junctions level illustration of the contagion traffic flow model. In the mean field limit, the aggregation of inflow and outflow intersections like those in the Figure, give a coarse grained description equivalent to a Susceptible-Infected-Susceptible (SIS) contagion process under stationary flow conditions. If the total transport is constrained to occur in a given time interval, the mean field model is then equivalent to a Susceptible-Infected-Recovered (SIR) process.
Urban traffic has been shown to be consistent with an out of equilibrium granular flow process [nobottle]. Consider therefore a network of links (edges), each representing idealized roads through which such a flow occurs. It’s additionally assumed that aggregates of cars traversing a link $n$ can be described in terms of a continuous car density flow in any point of the link except at the crossroads, which can be regarded as contact points in which vehicles enter or leave the $n$-th link. The master equation, or instantaneous transition time limit of the Chapman-Kolmogorov equation [VanKampen], for the car density at an edge $n$ is given by,
$\frac{d\rho_{n}(t)}{dt}=\sum_{m\neq n}q_{t}(n|m)\rho_{m}(t)-\sum_{r\neq n}q_{t}(r|n)\rho_{n}(t)$ (1)
The $q_{t}(n|m)$ and $q_{t}(r|n)$ terms in Eq.(1) represent the instantaneous flow rates of cars from the edges $m$ and $n$ to their directly connected edges $n$ and $r$, respectively. The precise description of $q_{t}(n|m)$ and $q_{t}(r|n)$ involve nonlinear interactions of vehicles at both sides of a given juncture due to the compromise between the vehicles in both lanes. In this sense, traffic signaling can be understood like a regulator of the inherent instabilities at street crossroads. Consider the flows, densities and average speeds in the vicinity of the intersections $n\leftarrow m$, where the arrow indicates transport from $m$ to $n$, $m\neq n$ and at $r\leftarrow n$, $r\neq n$ which vent transport outside the $n$-th road. In the present work we propose an average description of the traffic network junctions such that pockets of vehicles with densities $\rho_{k\leftarrow u}$ are transmitted in an intersection $k\leftarrow u$ with an average rate of $\bar{v}_{k\leftarrow u}$, so
$q_{t}(n|m)=\bar{v}_{n\leftarrow m}\rho_{n\leftarrow m}(t)$ (2)
$q_{t}(r|n)=\bar{v}_{r\leftarrow n}\rho_{r\leftarrow n}(t)$
The Eq. (2) can be interpreted like an extension of the definitional relationship between flow and density [libre, blubook-vol1-v10], $q=\bar{v}\rho$. Our proposed description broadens the definitional relationship by considering the $\bar{v}_{k\leftarrow u}$ terms to be the transport rates of densely packaged groups of vehicles at the neighborhood of a given intersection $k\leftarrow u$. Substituting the Eq. (2) in the master equation Eq. (1),
$\frac{d\langle\rho_{n}\rangle}{dt}=\sum_{m\neq n}\langle\bar{v}_{n\leftarrow m}\rho_{n\leftarrow m}\rho_{m}\rangle-\sum_{r\neq n}\langle\bar{v}_{r\leftarrow n}\rho_{r\leftarrow n}\rho_{n}\rangle$ (3)
The brackets in Eq. (3) denote spatial averages over the joint density of vehicles traversing the network. In the zeroth order approximation of vanishing vehicle density fluctuations, the mean field dynamics reads,
$\frac{d\langle\rho_{n}\rangle}{dt}=\langle\rho_{n}\rangle\left[\sum_{m\neq n}\langle\bar{v}_{n\leftarrow m}\rangle\langle\rho_{m}\rangle-\sum_{r\neq n}\langle\bar{v}_{r\leftarrow n}\rangle\langle\rho_{r}\rangle\right]$ (4)
The last term inside the square brackets in Eq. (4) is simply the venting rate of vehicles outside the $n$-th link. The first term inside the brackets gives the total vehicle transmission of incoming vehicles from links other than $n$. By normalizing spatially over the subgraph of links directly connected to $n$, the following coarse grained dynamics is obtained,
$\frac{d\rho_{n}(t)}{dt}=\beta_{n}\rho_{n}(1-\rho_{n})-\gamma_{n}\rho_{n}$ (5)
The mean field expression Eq. (5) is equivalent to a Susceptible-Infected-Susceptible (SIS) contagion process by identifying $\beta_{n}$ and $\gamma_{n}$ like contagion and recovery rate parameters, respectively. In the present context, these refer to the effective transmission and venting of vehicles to a link $n$ by its connected network of inwards and outwards links. If a total number of vehicles $C_{n}$ traverse the $n$-th link during a time period $T$, then under conservative conditions the density $\rho_{n}(t)$ should be normalized in time as,
$\frac{d\rho_{n}}{dt}=\rho_{n}\left[(1-\rho_{n})\beta_{n}-\gamma_{n}\right]\quad\text{subject to}$ (6)
$\int_{0}^{T}\rho_{n}(t)dt=C_{n}$
The constraints in Eq. (6) can be satisfied by introducing the temporal normalization factors,
$I_{n}\equiv\int_{0}^{T}\rho_{n}(t)dt,$ (7)
$S_{n}\equiv\int_{0}^{T}(C_{n}-\rho_{n})dt\equiv\int_{0}^{T}s_{n}(t)dt$
$R_{n}\equiv C_{n}-S_{n}-I_{n}$
The Eq. (7) states that the transport that traverse the $n$-th link during a time interval $T$ is conserved. The introduction of the constraints Eq.(6) therefore lead to a Susceptible-Infected-Recovered (SIR) type model for the traffic flow over the $n$-th node, in which a time span $T$ with total transport $C_{n}$ can be interpreted as an individual contagion process. By combining with the vanishing spatial density fluctuations limit of Eq. (5), the following coarse grained SIR description of an individual link in a mean field network is obtained,
$s(t)$ $=$ $s(0)e^{-\beta\rho(t)}$ (8)
$\frac{d\rho(t)}{dt}$ $=$ $\beta\rho(t)s(t)-\gamma\rho(t)$
2.1 Comparison with related approaches and the Fundamental Diagram of Traffic
Descriptions of traffic based on contagion dynamics models have been previously proposed, but these depend on ad-hoc definitions of congested states. For instance, in the seminal work [sun], congestion is defined by mapping to bond percolation on a small world network, where the congested states can be transmitted according to a contagion rate parameter and a percolation probability parameter. More recently in [saberi], congested states are formalized in terms of a threshold for the instantaneous average speed on a link relative to its maximum speed. Given this definition for a congested state road, is then proposed in [saberi] that congestion spreads through the network by a mechanism that at the aggregated level is described by a simple compartmental SIR dynamics. Ample empirical evidence is given in [saberi], that indicates that the compartmental SIR behavior closely match the temporal evolution of congestion in a given city.
In our model, congestion is not treated like a state of a link or group of links, but rather emerges in the form of a shape for flow vs density given a total transport in a given time. From Eq. (8) it follows that (with $s_{0}\equiv s(0)$), $\dot{\rho}(\rho)$ is expressed by,
$\dot{\rho}(\rho)\approx\beta\rho s_{0}e^{-\beta\rho}-\gamma\rho$ (9)
The resulting shape is consistent with the Fundamental Diagram of Traffic (FDT) [blubook-vol1-v10], in the sense that it permits to distinguish between free flow and congested phases by the sign of the derivative of the flow with respect to the density.
From Eq. (9), it follows that the macroscopic critical density point is approximately given by the solution of the transcendental equation,
$\rho_{c}=1-\frac{\gamma}{\beta s_{0}}e^{\beta\rho_{c}}$ (10)
The result Eq. (10) clearly expresses the value of our mean field approach for the understanding of traffic in terms of contagion dynamics. Previous epidemic models of road networks, like the one presented in [saberi], are largely reactive: they track the spread of a jam after certain speed threshold is breached. Our result Eq. (10) in contrast, offers a predictive capability. It implies that the “tipping point” into congestion is not a universal constant but is a function of the inflow or infection rate ($\beta$), the outflow or recovery rate ($\gamma$) and $s_{0}$ (the network load). This suggests that a traffic controller could theoretically increase the critical density (making the road more resistant to jams) by altering $\gamma$ (e.g., adjusting traffic light cycles to increase venting) relative to $\beta$. This provides a theoretical basis for “immunizing” intersections against jams by tuning their topological parameters, rather than just managing queues once they form.
3 Comparisons with empirical data
3.1 Direct experimental evidence
To have data for the direct empirical evidence of the proposed mean field model at the single link level requires precise flow measurements from all the relevant intersections in an entire subgraph. This can be difficult due to the usually incomplete coverage provided by traffic sensors. Here however we present an empirical study that intends to overcome the stated obstacle by analyzing vehicle flow data in a lane segment close to an intersection where a group of important traffic links meet. The case study is situated in the city of Saltillo, capital of the state of Coahuila in Mexico. The data has been gently provided by Autonomous University of Coahuila professor Dr. Jaime Burgos García together with the Municipal Institute of Sustainable Urban Mobility of Saltillo. The experimental setup has been implemented as follows. A vehicle count station for the road segment of interest, which is part of the Venustiano Carranza Boulevard in Saltillo, has been situated at the approximate latitude and longitude coordinates (25.459628,100.983199)(25.459628,-100.983199). Another vehicle count station has been placed nearly $1.0$ Km in direction north, at the approximate coordinates (25.466904,100.979632)(25.466904,-100.979632). The second station monitors the already mentioned hub of lanes, one of those corresponding to the chosen road section of the Venustiano Carranza Boulevard. By the combined vehicle count measurements of both stations, which haven taken in $15$ minutes time intervals, the number of cars entering and leaving the road segment through the hub and the total number of vehicles in the segment have been registered. The aggregated data is presented in the Table 1, in which the number of vehicles traversing the lane segment at the hub intersection is calculated as the difference between the total input plus output vehicles and the number of vehicles incoming from the Venustiano Carranza Boulevard road segment itself and continuing their transit along it. A total of $20$ measurements have been made the $25$th of September, $2024$. The first ten measurements have been taken at the morning traffic peak, which goes from $7:00$ AM to $9:30$ AM, while the second half of the dataset instances correspond to the afternoon peak hour, from $17:00$ PM to $19:30$ PM, local times.
We have used the provided data to test our mean field contagion model by normalizing the vehicles counts per measurement considering the total transport through the intersection during the total observation time span. Therefore the car density flux at the intersection is normalized as,
$\Delta\rho/\Delta t=\frac{(\text{number of vehicles traversing the intersection})}{(20*\text{total transport})}$ (11)
In this way, the data in the last column of the Table 1 gives a vehicle flux time series that we have used to fit the discrete Euler approximation of the SIR model, Eq.(8), neglecting fluctuations. The results are shown in the Figure 3, where the dotted lines represent the flow data, the dashed blue lines its 9090% confidence interval and the middle dashed red line gives the fitted SIR model. The predicted average rate parameters turn out to be $\beta=24.435\pm 4.0$, $\gamma=23.205\pm 4.0$, which are consistent with the actual observed rates of $25$ and $27$ vehicles per minute, respectively. Figure 4 on the other hand, uses the predicted mean $\beta$ and $\gamma$ rates to evaluate the function $\dot{\rho}(\rho)$, which gives the Fundamental Diagram of Traffic (FDT) for the considered road segment at the considered time spans. The large confidence intervals shown in Figure 3 confirm that significant fluctuations are present, which naturally leads to scatter in the FDT representation of Figure 4. Besides the statistical noise resulting from the limited sample size, the empirical data also displays the phenomenon of hysteresis loop in the Fundamental Diagram, which is considered to be a consequence of drivers behavior during acceleration and deceleration [traffhys]. It’s interesting to notice that our model successfully identifies the attractor (the red line) around which the real, noisy, hysteretic data (the black line) oscillate. That is, our mean field model successfully captures the central tendency (the “backbone”) of the Fundamental Diagram, effectively averaging through microscopic interactions like those resulting in the hysteresis loop of the experimental data.
Table 1: Vehicles count data at the peak hours in the vicinity of a highways intersection in Saltillo, Mexico
15 minutes intervals Input vehicles Output vehicles Vehicles in the Vehicles traversing
segment of link the intersection
1 466 482 418 530
2 502 461 422 541
3 444 435 372 507
4 521 486 427 580
5 445 413 344 514
6 428 422 359 491
7 364 384 307 441
8 454 424 363 515
9 373 373 317 429
10 256 295 223 328
11 294 369 228 435
12 373 461 275 559
13 356 466 270 552
14 334 446 268 512
15 253 370 182 441
16 221 303 162 362
17 204 218 131 291
18 284 323 230 377
19 343 415 295 463
20 443 466 353 556
Average rates per minute 25 27
Total transport 9424
Figure 3: Fitting of the SIR mean field model to experimental data in the city of Saltillo, Mexico (see Subsection 3.1 for a thorough explanation).
Figure 4: Experimental vehicle flow vs density data from the city of Saltillo, Mexico (black dotted lines) and the FDT function that results from the estimated SIR model (red dotted dashed lines).
3.2 Comparisons with traffic sensors data
Our contagion network traffic flow model has also been tested on representative instances of the publicly available UTD19 database [loder]. This database provides fundamental traffic variables including flow and density, from 41 cities worldwide. The data is collected from stationary traffic sensors, most of them inductive loop detectors, which report traffic flow and the fraction of time during an observation period in which a detector is occupied as a proxy for the vehicle density.
The upper and lower panels of the Figure 5, show the fitting of the SIR density dynamics from a single detector over a particular street, in the Aussersihl district of Zurich, Switzerland, over a $24$ hour period. This instance of the UTD19 database has been showcased in a number of works and has good quality sensor data [loder]. Fitting of the SIR rate parameters is done by minimizing the total squared error between the observed densities and those predicted by the SIR dynamics integrated by first order discretization through the Euler method. Because the street density data is a proxy, additional parameters for a linear transformation that resolves the particular sensor calibration are introduced. The resulting optimization problem has two levels. Firstly the rate parameters are estimated by solving
$\textit{min}_{\beta_{n},\gamma_{n}}\quad\sum_{t=1}^{T}[\hat{\rho}_{n}(\beta_{n},\gamma_{n},t)-\rho_{n}(t)]^{2}$ (12)
In Eq.(12), $\hat{\rho}_{n}$ is the vehicle density in the $n$-th link given by the one step ahead prediction of the SIR model at the observed time $t$. The sum of squared errors is calculated from the total number of observations $T$, over the $24$ hour period. The local fitting of the SIR dynamics has been done using the Broyden–Fletcher–Goldfarb–Shanno algorithm, assuming time independent rate parameters, $\beta_{n}(t)=\beta_{n}$, $\gamma_{n}(t)=\gamma_{n}$ and also assuming that $s(0)=1$.
In a second stage, optimal calibration parameters $s_{0},c_{1},c_{2}$ are calculated by,
$\textit{min}_{s_{0},c_{1},c_{2}}\quad\sum_{t=1}^{T}\left[\hat{y}_{n}(s_{0},c_{1},c_{2},t)-\frac{d\rho_{n}(t)}{dt}\right]^{2}$ (13)
$\hat{y}_{n}(s_{0},c_{1},c_{2},t)\equiv s_{0}\beta_{n}\rho_{n}(t)e^{-\beta_{n}\rho_{n}(t)}-\gamma_{n}\rho_{n}(t)$
The calibration parameter $0<s_{0}<1$ estimates the fraction of the overall transport covered by the $n$-th link in the mean field limit. The parameters $c_{1}$ and $c_{2}$ are mere calibration constants for the rates, $\beta_{n}\to c_{1}\beta_{n}$, $\gamma_{n}\to c_{2}\gamma_{n}$.
The upper panel of the Figure 5 displays the estimated and observed FDT, $\hat{y}_{n}(\rho_{n})$ and $\dot{\rho}_{n}(\rho_{n})$ respectively. The lower panel on the other hand, shows the estimated and observed SIR temporal dynamics of the vehicle density. Figure 6 reports the comparison between the predicted and observed total transport in the $24$ hour period for an arbitrary set of detectors from links in the vicinity and same district of the detector reported in the Figure 6. The links rate parameters are estimated by the previously discussed procedure individually. The estimated rate parameters averaged over the group of links are consistent with what expected from the equation (8).
Figure 5: Traffic dynamics from a single detector of the Schimmelstrasse street in the Aussersihl district of Zurich, Switzerland, at Wednesday 28th October, 2015. Empirical data is shown in black while the fit of the local mean field contagion model to the data is shown in red.
Figure 6: Predicted (red triangles) and observed (black circles) transport in 22 detectors from the Aussersihl district in Zurich, Switzerland, at Wednesday 28th October of 2015. From the estimated transmission rate parameters, the mean field crossroad contagion network traffic flow model predicts an average inflow over the covered region of $30.92$ veh/hr and an average venting rate of $6.54$ veh/hr. The arbitraily chosen detectors from the UTD19 database are, K14D15, K14D13, K14D14, K17D13, K17D14, K17D15, K17D12, K17D11, K13D14, K13D13, K13D12, K14D12, K11D19, K11D11, K12D16, K133D11, K133D12, K133D13, K11D12, K11D14, K12D14, K12D17.
4 Discussion
Despite being an idealization, the contagion framework to traffic appears to capture features that are essential to characterize urban transport at different spatial and temporal scales. The formalism integrates granular with continuous aspects of vehicle flow in conjunction with basic temporal constraints, leading to a local mean field contagion type model. Our empirical comparisons indicate that single detector data aggregates information from all of the links directly connected to the detector’s road. In fact, the experiment presented in Subsection 3.1 indicates that it’s possible to estimate macroscopic traffic network parameters, like total transport or optimal vehicle transit rates over a region, from purely local flow measurements at an intersection.
The macroscopic mean field equation (8), predicts that aggregating links in a given macroscopic area of the roads network should converge to macroscopic rate parameters that reflect the vehicle flow in the corresponding network’s section. This appears to be consistent with the empirical study reported in the Figure 6.
Clearly a next important step in the development of the proposed mean field theory, is to apply our approach to large macroscopic traffic networks datasets, by carefully taking into account the network topology. Our equation (10) indicates that the free flow optimal state of a macroscopic traffic network occurs in a critical density given in terms of aggregated rate parameters $\beta$, $\gamma$ and a normalization parameter associated with the total transport in the given time period, $s_{0}$. This may lead to “immunization” traffic control strategies, as explained in Subsection 2.1.
Generalizations to the mean field setup that consider network’s spatiotemporal structures in greater detail, can in principle be explored by taking advantage of recent developments in the study of contagion models on networks with complex topologies, like for instance those presented in [palafox].
5 Acknowledgments
The authors are grateful with Dr. Jaime Burgos García and Dr. Simón Rodríguez Rodríguez for the very useful discussions with them and particularly to Dr. Burgos García for having kindly shared with us the experimental data from the Municipal Institute of Sustainable Urban Mobility of Saltillo analyzed by us in Subsection 3.1.
The authors also acknowledge to the creators and maintainers of the UTD19 database, from which the traffic sensors data used in Subsection 3.2 has been taken. The UTD19 database is hosted in the site utd19.ethz.ch